You have a chance to buy an annuity that pays $3,300 at the beginning of each year for 3 years. You could earn 5.5% on your money in other investments with equal risk. What is the most you should pay for the annuity?
Answer #1

3300 3300 3300 نيا Present value + 3300 = 3300 + 3300 1.0552 160551 = 3300 + 3127.96 + 2964.89 = $9392.86 The most that we sh3300 3300 3300 نيا Present value + 3300 = 3300 + 3300 1.0552 160551 = 3300 + 3127.96 + 2964.89 = $9392.86 The most that we sh